OpenStudy (saadi101):

HELP! :P Solve using elimination: -x-7y=14 and -4x-14y=28

OpenStudy (welshfella):

multiply the first equation by -4 then add the result to the second eqution this will wliminate the x term

OpenStudy (welshfella):

-x-7y=14 * -4 becomes 4x + 28y = -56 now add this to -4x - 14y = 28

OpenStudy (welshfella):

note that 4x + (-4x) = 0

OpenStudy (welshfella):

now add the y terms and the numbers to get an equation only in y terms

OpenStudy (saadi101):

This is still going over my head a bit :/